OpenStudy (anonymous):

Please Help Gets Medal If Answers!!!!!!!!!!!!!!!! Max was solving the equation below and isn’t sure if his answer is correct. Explain to Max how he can check his answer and then help him identify any errors he made. Provide the correct solution in your explanation. 3x+6 8 = 7x−1 6 24 1 ( 3x+6 8 ) = 24 1 ( 7x−1 6 ) 3(3x+6) = 4(7x−1) 9x+18 = 28x−1 9x−9x+18 = 28x−9x−1 18 = 19x−1 18+1 = 19x−1+1 19 = 19x 1 = x

OpenStudy (anonymous):

I don't get it

OpenStudy (anonymous):

May you put it in a way I can understand it please.

OpenStudy (anonymous):

3x+6 8=7x−1 6 24 1 (3x+68)=241 (7x−16) this is the only way i can do it sorry

OpenStudy (anonymous):

what is the 6 24 1 do * or - or +

OpenStudy (anonymous):

oh ok I think i got a way to fix it and put it better

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

...

OpenStudy (anonymous):

@Compassionate (I am sorry bro but I have to go so I am calling a friend.

OpenStudy (anonymous):

3x+6 7x-1 ----- = ------ 8 6 24(3x+6) 24( 7x-1) ---- ---- = ---------- 1 ( 8 ) 1 ( 6 ) 3(3x+6) = 4(7x-1) 9x + 18 = 28x-1 9x-9x+18 = 28x-9x-1 18 = 19x-1 18+1 = 19x-1+1 19 = 19x 1 = x

OpenStudy (anonymous):

sory i took so long

OpenStudy (anonymous):

can someone help me please!!!!!!

OpenStudy (anonymous):

wait, gimme a sec, i'm trying to understand the equation

OpenStudy (anonymous):

k

OpenStudy (anonymous):

3(3x+6) = 4(7x-1) 9x + 18 = 28x-1 ---> should be 28x-4

OpenStudy (anonymous):

otherwise, your solution is okay

OpenStudy (anonymous):

im back my bad i went to do something

OpenStudy (anonymous):

@Chibi_Robo3

OpenStudy (anonymous):

@kjones0331